OpenStudy (anonymous):

Need help! Thank you. Will fan + medal! 1. Does the equation represent a direct variation> If so, find the constant variation. 5x = -3y (1 point) a. no b. yes; k = -5/3 c. yes; k = 5/3 d. yes; k = -3/5 2. Suppose y varies with x, and y = 10 when x=-3. What direct variation equation relates x and y? What is the value of y when x = -1? (1 pony) a. y = -3/10 x; 3/10 b. y = 1/10 x; -3/10 c. y = 10/3 x; -10/3 d. y = -10/3 x; 10/3 3. Suppose y varies directly with x, and y = 8 when x = -6. What direct variation equation relates x and y? What is the value of y when x = -2? (1 po

OpenStudy (anonymous):

Need help! Thank you. 1. Does the equation represent a direct variation> If so, find the constant variation. 5x = -3y (1 point) a. no b. yes; k = -5/3 c. yes; k = 5/3 d. yes; k = -3/5 2. Suppose y varies with x, and y = 10 when x=-3. What direct variation equation relates x and y? What is the value of y when x = -1? (1 pony) a. y = -3/10 x; 3/10 b. y = 1/10 x; -3/10 c. y = 10/3 x; -10/3 d. y = -10/3 x; 10/3 3. Suppose y varies directly with x, and y = 8 when x = -6. What direct variation equation relates x and y? What is the value of y when x = -2? (1 point) a. y = -0.75x; 1.50 b. y = -1.33xx; 2.67 c. y = 1.33x; -2.67 d. y = 0.13x; -0.25 4. The equation of the line on the graph below is is a direct variation equation. What is the constant variation? Need help! Thank you. 1. Does the equation represent a direct variation> If so, find the constant variation. 5x = -3y (1 point) a. no b. yes; k = -5/3 c. yes; k = 5/3 d. yes; k = -3/5 2. Suppose y varies with x, and y = 10 when x=-3. What direct variation equation relates x and y? What is the value of y when x = -1? (1 pony) a. y = -3/10 x; 3/10 b. y = 1/10 x; -3/10 c. y = 10/3 x; -10/3 d. y = -10/3 x; 10/3 3. Suppose y varies directly with x, and y = 8 when x = -6. What direct variation equation relates x and y? What is the value of y when x = -2? (1 point) a. y = -0.75x; 1.50 b. y = -1.33xx; 2.67 c. y = 1.33x; -2.67 d. y = 0.13x; -0.25 4. The equation of the line on the graph below is is a direct variation equation. What is the constant variation?  a. 1/4 b. 1/2 c. 3/4 d. 1 5. For the data in the table, does y vary directly with x? If it does, choose an equation for the direct variation. x y 8 11 16 22 24 33 a. yes; y = 2.75x b. yes; y = 0.6875x c. yes; y = 1.375x d. no; y does not vary directly with x